OpenStudy (anonymous):

Can someone please help me?? D: Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

First you'll need to convert the word problem to mathematical notation. What do you think it is?

OpenStudy (anonymous):

3sqrt(2x) + xsqrt(8x) - 5sqrt(18x

OpenStudy (anonymous):

Here's the formula for what you just typed (easier to read this way): \[3\sqrt(2x) + x \sqrt(8x) - 5\sqrt(18x)\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

-2sqrt(2x) .. is this right?

OpenStudy (anonymous):

Don't believe so. Show your work. Hint: Start out by figuring if there's a way to simplify any of the items that are being rooted (e.g. 8x and 18x). Note: \[\sqrt{ab} = \sqrt{a}\sqrt{b}\]